Let z1 and z2 be nth roots of unity which subtend a right angle at origin ,then prove that n must be of the form 4K

Grade:

1 Answers

Vikas TU
4 years ago
Dear student 
Zn -1 = 0 
Z = 11/n
Z1 = et2*pie/n 
Z2 = et2*pie/n 
Q1 = 2*pie*k / n 
Q2 = 2*pie(k+1)/n 
Q2 – Q1 = pie/2
2pie/n (K+1-k) = pie/2
n =4k